OpenStudy (anonymous):

Which equation represents the solution to the following? I have some lemons. Jill has two more limes than five times the number of lemons that I have. Emma has one-third as many limes as Jill. Emma has 4 limes. How many lemons do I have?

OpenStudy (anonymous):

@aric200 sorry I couldnt repost it yesterday :D

OpenStudy (aric200):

again?

OpenStudy (aric200):

I thought we established the answer was 2

OpenStudy (anonymous):

But noooo i need the equation v-v

OpenStudy (aric200):

... I waited 10 hours

OpenStudy (anonymous):

Im sorry I got yelled atttt

OpenStudy (aric200):

For 10 hours?? haha you are fine... Do i get multiple choices?

OpenStudy (anonymous):

I got yelled at and wasnt allowed on for 4 hours, and was unable to get online for another 6 :3 But ya. I can tell you the answer choices?

OpenStudy (aric200):

Please do

OpenStudy (anonymous):

\[ 5L+2=\frac{ 4 }{ 3 } , \frac{ 5L+2 }{ 3 } =4 , \frac{ 5L }{ 3} +2=4 , \frac{ 5 }{ L } + \frac{ 1 }{ 3}+2=4\]

OpenStudy (anonymous):

Sorry if thats confusing o.o

OpenStudy (abhisar):

Let x be the number of lemon i have, then \(\huge\frac{2+5x}{3}\) = 4

OpenStudy (abhisar):

can u solve it now ?

OpenStudy (anonymous):

The answer is two? I thought I already got the number of lemons I had(or Aric did) but ya? Im confused :I

OpenStudy (abhisar):

yes the answer will be 2. Jill will have 12 and emma will have 4

OpenStudy (abhisar):

u want me to xplain ?

OpenStudy (anonymous):

I dont need it explained, Aric got 2 yesterday and they helped explain how they got it. I kind of only needed the equation

OpenStudy (abhisar):

ok

OpenStudy (anonymous):

Thanks for the help :)

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)